OpenStudy (plasmataco):

I just gotta make sure of something. The derivative and intergral of e^x is still e^x right?

OpenStudy (plasmataco):

@Vocaloid @Rushwr @isaac4321

OpenStudy (freckles):

yes \[\frac{d}{dx}e^{x}=e^x \text{ and } \int\limits e^x dx=e^x+C\]

OpenStudy (plasmataco):

Oh. Ok :3 thx

OpenStudy (plasmataco):

So... The intergral of e^kx is equal to e^kx?

OpenStudy (plasmataco):

Or is it different

ganeshie8 (ganeshie8):

with respect to k or x ?

OpenStudy (freckles):

\[\int\limits_{}^{}e^{kx} dx \\ u=kx \\ du=k dx \\ \int\limits e^{u} \frac{1}{k} du =\frac{1}{k} \int\limits e^{u} du=...\]

OpenStudy (plasmataco):

With respect of x I think.

OpenStudy (plasmataco):

One more :3

OpenStudy (plasmataco):

What's the derivitave of ln(x)

OpenStudy (freckles):

\[y=\ln(x) \\ e^{y} =x \] you know how to differentiate e^y w.r.t x right?

OpenStudy (plasmataco):

No

OpenStudy (freckles):

apply chain rule

OpenStudy (plasmataco):

Ok...

OpenStudy (plasmataco):

Ohhhh so would it be 1/x?

OpenStudy (freckles):

\[\frac{d}{dx}e^x=e^x \\ \frac{d}{dx}e^{u}=u' e^{u} \\ \frac{d}{dx}e^{y}=y'e^{y} \\ y=\ln(x) \text{ is equivalent \to } e^{y}=x \\ \text{ differentiating both sides gives } y'e^{y}=1 \\ y'=\frac{1}{e^y}=\frac{1}{x} \\ \text{ since } e^{y}=x\]

OpenStudy (plasmataco):

Or not...?

OpenStudy (freckles):

you are right

OpenStudy (plasmataco):

Oh.. Yay thx @freckles helps a lot

OpenStudy (freckles):

I was just going off by what we had earlier I think the definition of ln(x) is... \[\ln(x)=\int\limits_1^x \frac{1}{t} dt\]

OpenStudy (freckles):

and the you use the fundamental theorem of calculus to find the derivative of ln(x) w.r.t. x

OpenStudy (freckles):

where x>0